# ZERO: Industry-ready Vision Foundation Model with Multi-modal Prompts

### v3

@@ 1 Introduction
Historically, improving model performance required deep domain expertise to inject task-specific inductive biases into model architectures. However, the advent of large-scale datasets has shifted this paradigm. Empirical evidence increasingly supports that stacking generic Transformer layers, rather than hand-crafting domain-specific modules, can yield superior results [3]. This shift has simplified model design and reduced reliance on manual heuristics.
This transition has given rise to foundation models, which have substantially accelerated progress across a wide range of AI tasks. Text-centric models like BERT [2], vision-centric models such as DINOv2 [24], and multi-modal models including CLIP [25] have demonstrated remarkable generalization capabilities. These advances have been fueled by the release of massive public datasets (e.g., SA-1B [15], LAION-5B [30]) and parameter-efficient fine-tuning methods like LoRA [8], which enable fast adaptation with minimal computational cost.
Despite these advancements, current foundation models remain ill-suited for zero-shot deployment in real-world, domain-specific industrial settings. The datasets used to train them largely focus on general-purpose content, while high-quality, domain-specific datasets, especially in areas such as manufacturing, logistics, and medical imaging, are scarce due to high annotation costs and long-tail data distributions. Generative models can support data augmentation, but their benefits are often constrained by diminishing returns, as predicted by scaling laws [37, 4]. Recent work on active learning, data valuation, and subset selection [6, 32] aims to mitigate these issues, but the challenge of curating effective domain-specific datasets persists as a core bottleneck to industrial AI adoption.
@@ 2.1 Open-vocabulary object detection
Recent advancements in contrastive learning have enabled models to relate images and text by mapping text and image embeddings into a shared semantic space [25, 38, 31]. This breakthrough has allowed open-vocabulary object detectors to be trained successfully on large-scale web-based image datasets without extensive relabeling.
Pioneering works, including OWL-ViT [21] and GLIP [17], used contrastive learning methods to align image and text embeddings. GroundingDINO [18] significantly improved performance by integrating modern object detection architectures with language models in a dual-encoder framework. GroundingDINO became the foundation for subsequent research, including MM-GroundingDINO [40], which extended capabilities to multi-modal inputs, and Grounded-SAM [28], which combined SAM [26] for segmentation. Most recently, LLM-Det [5] has advanced this research direction by incorporating large language models, enabling models to understand more sophisticated natural language queries.
@@ 2.2 Object detection with visual prompts
Another line of research extends open-vocabulary object detection by incorporating visual prompts to alleviate the inherent ambiguity in text prompts. This approach includes DINOv [16], T-Rex [9], and T-Rex2 [10]. DINOv, built on top of DINO [39], enables visual in-context prompting, where visual prompts come from the same input images in which the model searches for objects. T-Rex and T-Rex2 relax this constraint of in-context prompting, allowing models to use visual examples from any image source. DINO-X[27] further pushed the boundaries by enabling models to present object masks and keypoints. It also introduced universal prompting, where the model locates all objects without user prompts.
@@ 2.3 Real-time visual grounding
One problem with previous models, despite being powerful, is that they require substantial memory and computational resources, which hinders practical deployment and real-time inference in industry. To address this problem, several works have implemented open-set object detectors based on YOLO architectures. YOLO-World [1] was first proposed for this purpose, and YOLOE [34] further improved performance in terms of mAP through more sophisticated architectural design. Although ZERO is not intended for real-time inference, we discuss our efficiency-oriented design choices for deployment in Section 3.2.2.
@@ 2.4 Referring expression comprehension
Recent improvements in vision-language models (VLMs) have opened up a new research direction involving more complex logical reasoning for object detection. Compared to previous works, this line of research focuses more on filtering detected objects using VLMs’ capabilities, enabling models to account for positional relationships between objects and object attributes. These works constructed their own datasets to train their models by integrating outputs from existing models and manual labeling into a loop. ChatRex [12] was the first attempt to merge VLMs with object detectors, and Rex-Seek [14] focuses more on processing human attribute queries. Rex-Thinker [13] further improved Rex-Seek by adding reasoning capabilities to the model.
@@ 3.1.3 Data labeling
The pipeline integrates two complementary directions. First, we extract captions from images using InternVL3 [41] and identify noun phrases via SpaCy [7]. These noun phrases are filtered using WordNet [20] to isolate expressions corresponding to physical objects. Using a preliminary version of ZERO trained on ground-truth annotations, we then aggregate inference results guided by both textual and visual prompts to localize bounding boxes for each phrase. Second, we reverse the process by generating object proposals with UPN and extracting corresponding captions using ChatRex [11]. In this step, the captioning focuses on visual attributes such as color, shape, and material, allowing for richer object-level descriptions. To maintain label quality, we filter out instances with extremely small bounding boxes and apply SigLIP2 [31] to assess alignment between cropped image regions and their generated captions. Low-alignment pairs are removed from the training set.
@@ 3.2 Training strategy
While numerous open-source models for the visual grounding task exist, they are typically limited to supporting a single type of prompt, most often textual. Some models like YOLOE [33] support visual prompting, but still underperform compared to modern open-vocabulary detectors [10, 27]. To overcome this limitation without redesigning the base model architecture, we focus on extending pretrained single-modal models to support multiple prompt modalities. In summary, our approach distills the open-vocabulary detection capabilities of a pretrained model in the early stages and subsequently enhances performance through multi-prompt synergy. The final trained model is efficient at inference time, requiring only the contrastive prompt encoders.
@@ 3.2.1 Progressive adaptation to multi-prompt
Naively appending a visual prompt encoder can lead to performance degradation, particularly during early training stages, due to the representational mismatch with the pretrained text prompt encoder. To mitigate this, we begin training with text-only prompts and employ an InfoNCE loss [23] to align the visual prompt encoder with the pretrained text prompt encoder. To prevent degradation of the text embeddings, we limit gradient propagation so that distillation happens only in one direction: from the text encoder to the visual encoder.
Relying solely on unidirectional distillation for training stability can limit the synergy between text and visual prompts. To address this, we incorporate a contrastive backbone, such as CLIP [25], which inherently supports text-visual alignment. In our framework, the visual prompt encoder is realized through the contrastive visual encoder, and a new contrastive text encoder is introduced to gradually replace the original pretrained text encoder.
@@ 4 Experiment
T-Rex2 [10] 6.5M 62.5 - 13.0 -
Florence-2 [36] 126M - 4.3 - -
YOLOE [33] 1.4M 50.5 7.5 12.0 15.0
OV-DINO [35] 2.4M - 10.2 - -
DINO-X [27] 20M $\sim$ 100M - 12.8 - -
OWLv2 [22] 2.1M - 12.9 - -
@@ 4.1 Zero-shot performance
We evaluate zero-shot detection performance under three prompting protocols introduced in Jiang et al. [10]. Text-G corresponds to the standard open-vocabulary setting, where object categories are specified using only text prompts (i.e., class names) without any visual exemplars. Visual-G employs category-level visual prompts derived by averaging the embeddings of $N$ training exemplars per class, simulating scenarios where a few reference images are available. Visual-I represents an interactive setting in which a single instance crop from the test image serves as the visual prompt to detect similar objects within the same image. While this assumes access to ground-truth regions at test time, it reflects practical applications such as interactive annotation. Results are summarized in Table 1. Note that cells marked with ‘–’ indicate that the corresponding prompt type is not supported for the given model.
@@ 4.1.1 Academic benchmark
ZERO demonstrates strong performance on the widely used LVIS-Val benchmark, achieving a Visual-I AP score of 60.1. This outperforms comparable open-source models such as YOLOE [33] (50.5) and approaches the state-of-the-art T-Rex2 [10] (62.5). Remarkably, ZERO attains this level of accuracy using only 0.9M training examples—significantly fewer than other models that rely on tens or even hundreds of millions of samples—highlighting its data efficiency and practical scalability.
@@ 4.2.1 Object Instance Detection challenge
From this baseline, we introduced a series of targeted improvements that reflect practical deployment constraints while maximizing performance under zero-shot conditions. First, category refinement was applied using a visual embedding model to align instance representations, combined with tiled inference to improve localization granularity. Next, to reduce spurious background predictions, a common issue in open-world detection, we trained a lightweight background classifier. Batched Non-Maximum Suppression (NMS) was then employed to handle overlapping predictions at scale efficiently. Finally, bounding box quality was enhanced using SAM2-based refinement [26]. The progressive enhancements are reported in Table 2 and demonstrate the effectiveness of modular inference-time improvements of ZERO that do not require additional training or fine-tuning.
@@ 4.2.2 Foundation Few-shot Object Detection challenge
Foundational FSOD challenge [19, 29] is designed to rigorously evaluate models’ capabilities in performing object detection with extremely limited annotations across diverse and heterogeneous industrial domains. Using the RF20VL-fsod dataset, which encompasses 20 distinct domains, participants are tasked with detecting objects given only 10 bounding box annotations per category. This sparse labeling setup realistically simulates practical constraints in industrial settings where exhaustive annotation is costly or infeasible.

### v4

@@ 1 Introduction
Historically, improving model performance required deep domain expertise to inject task-specific inductive biases into model architectures. However, the advent of large-scale datasets has shifted this paradigm. Empirical evidence increasingly supports that stacking generic Transformer layers, rather than hand-crafting domain-specific modules, can yield superior results [dosovitskiy2020image]. This shift has simplified model design and reduced reliance on manual heuristics.
This transition has given rise to foundation models, which have substantially accelerated progress across a wide range of AI tasks. Text-centric models like BERT [devlin2019bert], vision-centric models such as DINOv2 [oquab2023dinov2], and multi-modal models including CLIP [radford2021learning] have demonstrated remarkable generalization capabilities. These advances have been fueled by the release of massive public datasets (e.g., SA-1B [kirillov2023segment], LAION-5B [schuhmann2022laion]) and parameter-efficient fine-tuning methods like LoRA [hu2022lora], which enable fast adaptation with minimal computational cost.
Despite these advancements, current foundation models remain ill-suited for zero-shot deployment in real-world, domain-specific industrial settings. The datasets used to train them largely focus on general-purpose content, while high-quality, domain-specific datasets, especially in areas such as manufacturing, logistics, and medical imaging, are scarce due to high annotation costs and long-tail data distributions. Generative models can support data augmentation, but their benefits are often constrained by diminishing returns, as predicted by scaling laws [yamaguchi2023limitation, fan2024scaling]. Recent work on active learning, data valuation, and subset selection [go2023transferable, vo2024automatic] aims to mitigate these issues, but the challenge of curating effective domain-specific datasets persists as a core bottleneck to industrial AI adoption.
@@ 2.1 Open-vocabulary object detection
Recent advancements in contrastive learning have enabled models to relate images and text by mapping text and image embeddings into a shared semantic space [radford2021learning, zhai2023sigmoid, tschannen2025siglip]. This breakthrough has allowed open-vocabulary object detectors to be trained successfully on large-scale web-based image datasets without extensive relabeling.
Pioneering works, including OWL-ViT [minderer2022simple] and GLIP [li2022grounded], used contrastive learning methods to align image and text embeddings. GroundingDINO [liu2024grounding] significantly improved performance by integrating modern object detection architectures with language models in a dual-encoder framework. GroundingDINO became the foundation for subsequent research, including MM-GroundingDINO [zhao2024open], which extended capabilities to multi-modal inputs, and Grounded-SAM [ren2024grounded], which combined SAM [ravi2024sam] for segmentation. Most recently, LLM-Det [fu2025llmdet] has advanced this research direction by incorporating large language models, enabling models to understand more sophisticated natural language queries.
@@ 2.2 Object detection with visual prompts
Another line of research extends open-vocabulary object detection by incorporating visual prompts to alleviate the inherent ambiguity in text prompts. This approach includes DINOv [li2024visual], T-Rex [jiang2023trex], and T-Rex2 [jiang2024t]. DINOv, built on top of DINO [zhang2022dino], enables visual in-context prompting, where visual prompts come from the same input images in which the model searches for objects. T-Rex and T-Rex2 relax this constraint of in-context prompting, allowing models to use visual examples from any image source. DINO-X[ren2024dino] further pushed the boundaries by enabling models to present object masks and keypoints. It also introduced universal prompting, where the model locates all objects without user prompts.
@@ 2.3 Real-time visual grounding
One problem with previous models, despite being powerful, is that they require substantial memory and computational resources, which hinders practical deployment and real-time inference in industry. To address this problem, several works have implemented open-set object detectors based on YOLO architectures. YOLO-World [cheng2024yolo] was first proposed for this purpose, and YOLOE [wang2025yoloerealtimeseeing] further improved performance in terms of mAP through more sophisticated architectural design. Although ZERO is not intended for real-time inference, we discuss our efficiency-oriented design choices for deployment in Section 3.2.2.
@@ 2.4 Referring expression comprehension
Recent improvements in vision-language models (VLMs) have opened up a new research direction involving more complex logical reasoning for object detection. Compared to previous works, this line of research focuses more on filtering detected objects using VLMs’ capabilities, enabling models to account for positional relationships between objects and object attributes. These works constructed their own datasets to train their models by integrating outputs from existing models and manual labeling into a loop. ChatRex [jiang2024chatrextamingmultimodalllm] was the first attempt to merge VLMs with object detectors, and Rex-Seek [jiang2025referringperson] focuses more on processing human attribute queries. Rex-Thinker [jiang2025rexthinkergroundedobjectreferring] further improved Rex-Seek by adding reasoning capabilities to the model.
@@ 3.1.3 Data labeling
The pipeline integrates two complementary directions. First, we extract captions from images using InternVL3 [zhu2025internvl3] and identify noun phrases via SpaCy [Honnibal_spaCy_Industrial-strength_Natural_2020]. These noun phrases are filtered using WordNet [miller1995wordnet] to isolate expressions corresponding to physical objects. Using a preliminary version of ZERO trained on ground-truth annotations, we then aggregate inference results guided by both textual and visual prompts to localize bounding boxes for each phrase. Second, we reverse the process by generating object proposals with the automatic mask generation by SAM2 [ravi2024sam] and extracting corresponding captions using Qwen2.5-VL [Qwen2.5-VL]. In this step, the captioning focuses on visual attributes such as color, shape, and material, allowing for richer object-level descriptions. To maintain label quality, we filter out instances with extremely small bounding boxes and apply SigLIP2 [tschannen2025siglip] to assess alignment between cropped image regions and their generated captions. Low-alignment pairs are removed from the training set.
@@ 3.2 Training strategy
While numerous open-source models for the visual grounding task exist, they are typically limited to supporting a single type of prompt, most often textual. Some models like YOLOE [wang2025yoloe] support visual prompting, but still underperform compared to modern open-vocabulary detectors [jiang2024t, ren2024dino]. To overcome this limitation without redesigning the base model architecture, we focus on extending pretrained single-modal models to support multiple prompt modalities. In summary, our approach distills the open-vocabulary detection capabilities of a pretrained model in the early stages and subsequently enhances performance through multi-prompt synergy. The final trained model is efficient at inference time, requiring only the contrastive prompt encoders.
@@ 3.2.1 Progressive adaptation to multi-prompt
Naively appending a visual prompt encoder can lead to performance degradation, particularly during early training stages, due to the representational mismatch with the pretrained text prompt encoder. To mitigate this, we begin training with text-only prompts and employ an InfoNCE loss [oord2018representation] to align the visual prompt encoder with the pretrained text prompt encoder. To prevent degradation of the text embeddings, we limit gradient propagation so that distillation happens only in one direction: from the text encoder to the visual encoder.
Relying solely on unidirectional distillation for training stability can limit the synergy between text and visual prompts. To address this, we incorporate a contrastive backbone, such as CLIP [radford2021learning], which inherently supports text-visual alignment. In our framework, the visual prompt encoder is realized through the contrastive visual encoder, and a new contrastive text encoder is introduced to gradually replace the original pretrained text encoder.
@@ 4 Experiment
T-Rex2 [jiang2024t] 6.5M 62.5 - 13.0 -
Florence-2 [xiao2024florence] 126M - 4.3 - -
YOLOE [wang2025yoloe] 1.4M 50.5 7.5 12.0 15.0
OV-DINO [wang2024ov] 2.4M - 10.2 - -
DINO-X [ren2024dino] 20M $\sim$ 100M - 12.8 - -
OWLv2 [minderer2023scaling] 2.1M - 12.9 - -
@@ 4.1 Zero-shot performance
We evaluate zero-shot detection performance under three prompting protocols introduced in jiang2024t. Text-G corresponds to the standard open-vocabulary setting, where object categories are specified using only text prompts (i.e., class names) without any visual exemplars. Visual-G employs category-level visual prompts derived by averaging the embeddings of $N$ training exemplars per class, simulating scenarios where a few reference images are available. Visual-I represents an interactive setting in which a single instance crop from the test image serves as the visual prompt to detect similar objects within the same image. While this assumes access to ground-truth regions at test time, it reflects practical applications such as interactive annotation. Results are summarized in Table 1. Note that cells marked with ‘–’ indicate that the corresponding prompt type is not supported for the given model.
@@ 4.1.1 Academic benchmark
ZERO demonstrates strong performance on the widely used LVIS-Val benchmark, achieving a Visual-I AP score of 60.1. This outperforms comparable open-source models such as YOLOE [wang2025yoloe] (50.5) and approaches the state-of-the-art T-Rex2 [jiang2024t] (62.5). Remarkably, ZERO attains this level of accuracy using only 0.9M training examples—significantly fewer than other models that rely on tens or even hundreds of millions of samples—highlighting its data efficiency and practical scalability.
@@ 4.2.1 Object Instance Detection challenge
From this baseline, we introduced a series of targeted improvements that reflect practical deployment constraints while maximizing performance under zero-shot conditions. First, category refinement was applied using a visual embedding model to align instance representations, combined with tiled inference to improve localization granularity. Next, to reduce spurious background predictions, a common issue in open-world detection, we trained a lightweight background classifier. Batched Non-Maximum Suppression (NMS) was then employed to handle overlapping predictions at scale efficiently. Finally, bounding box quality was enhanced using SAM2-based refinement [ravi2024sam]. The progressive enhancements are reported in Table 2 and demonstrate the effectiveness of modular inference-time improvements of ZERO that do not require additional training or fine-tuning.
@@ 4.2.2 Foundation Few-shot Object Detection challenge
Foundational FSOD challenge [madan2024revisiting, robicheaux2025roboflow100] is designed to rigorously evaluate models’ capabilities in performing object detection with extremely limited annotations across diverse and heterogeneous industrial domains. Using the RF20VL-fsod dataset, which encompasses 20 distinct domains, participants are tasked with detecting objects given only 10 bounding box annotations per category. This sparse labeling setup realistically simulates practical constraints in industrial settings where exhaustive annotation is costly or infeasible.